In the laboratory a student finds that it takes 230 Joules to increase the temperature of 14.7 grams of gaseous nitrogen from 23.6 to 39.5 degrees Celsius. The specific heat of nitrogen calculated from her data is _ J/g°C.
Answer #1

Given information: q = 230 J m = 14.7g Ti = 23.6° C TE = 39.5°C AT = Tf - Ti = 139.5-23:6)°C = 15.9°CThe specific heat of nitrogen is calculated as follows: q = matcp cp =q maT = 230T - 14.78 x 15.9°C cp = 0.984 Ilgic